OpenStudy (anonymous):

The width of a rectangle is half as long as the length. The rectangle has an area of 18 square feet. What are the length and width of the rectangle?

OpenStudy (a_clan):

W=0.5L ............(1) W*L=18 sq ft ................(2) substitute (1) in (2) 0.5L*L=18 0.5(L^2)=18 L=6 ft W=3 ft

OpenStudy (anonymous):

Thank youuuuu :D

OpenStudy (anonymous):

A=L*W W=1/2L 18=L*(1/2L) 18=1/2L^2 \[L^2=18*\frac{2}{1}\] L^2=36 L=6